OpenStudy (anonymous):

A container holds 3 gallons of juice. How much is this in pints?

OpenStudy (studygurl14):

1 gallon = 4 quarts 1 quart = 4 cups 1 cup = 2 pints

OpenStudy (anonymous):

24 pints!

OpenStudy (studygurl14):

^That is incorrect

OpenStudy (igreen):

1 gallon = 8 pints, so what's 8 * 3? @

OpenStudy (igreen):

@iHomework

OpenStudy (igreen):

@StudyGurl14 How is that incorrect..she got it right..but she's not supposed to give the answer.

OpenStudy (studygurl14):

Wouldn't it be 3•4•4•2?

OpenStudy (studygurl14):

I'm confused..how are there 8 pints in a gallon, if the information i provided up there^ is true?

OpenStudy (anonymous):

I Home work needed help!

OpenStudy (anonymous):

IHomework*

OpenStudy (igreen):

No. 1 gallon = 8 pints. Therefore 3 gallons = 8 * 3 = 24 pints

OpenStudy (studygurl14):

Oh...sorry then.

OpenStudy (anonymous):

I am right.

OpenStudy (studygurl14):

But if there are 4 quarts in a gallon, 4 cups in a quart, and 2 pints in a cup, how is there only 8 pints in a gallon?

OpenStudy (igreen):

It's okay..everyone makes mistakes.

OpenStudy (igreen):

There are 1/2 pints in a cup..not 2.

OpenStudy (studygurl14):

Oh, I see my mistake! There are two cups in a pint, not 2 pints in a cup. Sorry everyone

OpenStudy (igreen):

So it would be: 4 * 4 * 1/2 = 8

OpenStudy (igreen):

You had the right idea!

OpenStudy (studygurl14):

I am so sorry...Okay, then 24 is most definitely the answer

OpenStudy (igreen):

Yep.